What are the dimensions of a flat screen T.V. with a surface area of 218.4 square inches?
Answer by Alan3354(69443)

What are the dimensions of a flat screen T.V. with a surface area of 218.4 square inches?
--------------
Most have an aspect ratio of 16:9, 16 wide by 9 high.
h = (9/16)w
-------
218.4 = w*h
218.4 = (9/16)w*w
w^2 = 218.4*16/9
w =~ 19.7 inches
h =~ 11.08 inches
------